Table 1: B-FIELD SIGNS FOLLOWING ARITHMETIC OPERATIONS

Unsigned  + Sign  − Sign
B field  B field  B field
if value is  if value is  if value is
A < B A = B A > B A < B A = B A > B A < B A = B A >
ADD UNSIGNED A NS  NS  NS  +  +  +  −  −0  +
ADD +A  NS  NS  NS  +  +  +  −  −0  +
ADD −A  +  +0  −  +  +0  −  −  −  −
SUB UNSIGNED A  +  +0  −  +  +0  −  −  −  −
SUB +A  +  +0  −  +  +0  −  −  −  −
SUB −A  NS  NS  NS  +  +  +  −  −0  +